The questions refer to a vertical cylindrical beaker holding water. The density of water is 1.0 g/cm3. The volume of a cylinder is πR2H, where R is the radius of the cylinder and H is the height.

A. If the beaker holds 1.88 kg of water and the bottom of the container has a diameter of 6.0 cm, what is the weight of the water in Newton's?

B. If the beaker holds 2.0 kg of water and the bottom of the container has a diameter of 6.4 cm, what is the height of the water centimeters?
Answer #1

A) We know, Weight = Mass * acceleration due to gravity

Weight = 1.88 * 9.8

Weight = 18.424 N

B) We know,

Density = \frac{Mass}{Volume}

Mass of water in grams = 2 * 1000 = 2000 gm

Radius of the cylinder = 3.2 cm

Volume of the cylinder = 3.14 * (3.2)2 * H, where H = height of water in centimeters

Putting values in the above equation:

1 = \frac{2000}{3.14*3.2^2*H}

2000=3.14*10.24*H

H = 62.2\ cm

PLEASE RATE IF YOU LIKE IT.